a rod is supported horizontally by means of two strings of equal length. of one of the string is cut . Then tension in other string at the same instant will..

Grade:11

2 Answers

Achal Sood
7 years ago
Initially, tension in both string is M/2. Let there be a point b on the end of the rod from which the string has been cut.Acc of b is l*alpha. Let  the end attached with string be a. Torque about a=moment of inertia of a *alpha.Now Mg*l/2=Ml2alpha/3. alpha=3g/l. ac=l/2alpha= 3g/4. Mg-T1=Mac. Mg-T1=M*3g/4. T1=Mg/4.
 
 
ankit singh
3 years ago

equationofforcegivesmgT=matorqueaboutthepointontherodwheretensionactsmg2l=Iα=3ml2αanda=α2lSolvingthesewegetT=mg43mg=41mginitiallyTwas21mg,sotheTensionisdecreased.